Doctor J. is considering purchasing a new blood analysis machine to test for HIV; it will cost $60,000. He estimates that he could charge $25.00 for an office visit to have a patient's blood analyzed, while the actual cost of a blood analysis would be $5.00. What would be his profit if he were to perform 5,000 HIV blood analyses? Question 13 options: $0 $40,000 $60,000 $25,000 $100,000
Answer #1

Fixed cost (FC) = $60000

Variable cost (VC) = $5

Revenue (R) = $25

Volume of output (Q) = 5000

Profit = Q(R-VC)-FC

= 5000(25-5)-60000

= (5000 × 20) - 60000

= 100000-60000

= 40000

So the answer is $40000